# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ 5 Conclusion
Generative LLMs like GPT series and Llama series are currently the main focus due to their high algorithm performance. The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. This paper presents a comprehensive survey of efficient generative LLM inference on different hardware platforms. We provide an overview of the algorithm architecture of mainstream generative LLMs and summarize different optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP. Furthermore, we perform a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms by considering hardware power consumption, absolute inference speed (tokens/s), and energy efficiency (tokens/J). And we also point to the future trends and potential developments of generative LLMs and hardware technology for edge-side scenarios.

### v3

@@ 6. Conclusion
Generative LLMs like GPT series and Llama series are currently the main focus due to their high algorithm performance. The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. This paper presents a comprehensive survey of efficient generative LLM inference on different hardware platforms. We provide an overview of the algorithm architecture of mainstream generative LLMs and summarize different optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP. Furthermore, we perform a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms by considering hardware power consumption, absolute inference speed (tokens/s), and energy efficiency (tokens/J). We point out that the development of edge intelligence has gained significant momentum, and three trends (multimodality, inference-time compute, and higher inference energy efficiency) are promising to redefine the capabilities of edge AI systems. Based on the software and hardware optimization, future edge chips are expected to achieve high throughput and low energy consumption, eliminating the 1-2 orders of magnitude gap in edge AI inference.